### v1

# Structuring the Unstructured: A Multi-Agent System for Extracting and Querying Financial KPIs and Guidance

Abstract.
Extracting structured and quantitative insights from unstructured financial filings is essential in investment research, yet remains time-consuming and resource-intensive. Conventional approaches in practice rely heavily on labor-intensive manual processes, limiting scalability and delaying the research workflow. In this paper, we propose an efficient and scalable method for accurately extracting quantitative insights from unstructured financial documents, leveraging a multi-agent system composed of large language models. Our proposed multi-agent system consists of two specialized agents: the Extraction Agent and the Text-to-SQL Agent. The Extraction Agent automatically identifies key performance indicators from unstructured financial text, standardizes their formats, and verifies their accuracy. On the other hand, the Text-to-SQL Agent generates executable SQL statements from natural language queries, allowing users to access structured data accurately without requiring familiarity with the database schema. Through experiments, we demonstrate that our proposed system effectively transforms unstructured text into structured data accurately and enables precise retrieval of key information. First, we demonstrate that our system achieves approximately 95% accuracy in transforming financial filings into structured data, matching the performance level typically attained by human annotators. Second, in a human evaluation of the retrieval task – where natural language queries are used to search information from structured data – 91% of the responses were rated as correct by human evaluators. In both evaluations, our system generalizes well across financial document types, consistently delivering reliable performance.
Financial Information Extraction, Multi-Agent Systems, Text-to-SQL, Financial Analytics
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/25/07conference: The 2nd Workshop on Financial Information Retrieval in the Era of Generative AI; July 17, 2025; Padua, Italy
1. Introduction
The financial industry generates vast volumes of unstructured documents-10-Ks, 10-Qs, 8-Ks, DEF14As, and investor memos-that are central to investment research and decision-making. Analysts have traditionally extracted key performance indicators (KPIs) from these documents by manually scanning text, copying metrics, normalizing formats, and writing internal summaries. This process is not only time-consuming but also error-prone and unscalable in modern research environments where timeliness and breadth of coverage are essential.
While recent large language models (LLMs) have shown promise in structuring unstructured data, their application in financial workflows remains unreliable. In the extraction phase, LLMs often misinterpret numeric guidance, extract semantically incorrect values (e.g., YoY instead of QoQ), or omit critical metrics entirely. These issues stem from a lack of domain-specific reasoning, layout ambiguity, and difficulty distinguishing factual signals from narrative commentary. Even when extraction is successful, querying this data via natural language introduces further challenges: text-to-SQL models are sensitive to phrasing, schema variation, and inherent LLM randomness, often leading to invalid or misleading responses.
These problems are compounded by the fact that most LLM-based systems rely on monolithic, single-shot inference pipelines that lack transparency, fail to isolate errors, and cannot iteratively refine intermediate outputs. This makes it difficult to enforce consistency, inject domain rules, or apply structured validation across stages.
To address these challenges, we propose a scalable, document-agnostic multi-agent framework for financial document understanding. Our system consists of two specialized agents: an Extraction Agent, which parses and validates KPIs using domain-tuned prompts and logic, and a Text-to-SQL Agent, which translates analyst queries into executable SQL over the structured output. By explicitly decomposing the workflow, the system allows targeted evaluation, modular tuning, and the integration of feedback loops at each stage-enabling iterative refinement and robust error handling.
This modular design not only automates manual workflows but also ensures both high-fidelity structuring and precise, trustworthy retrieval. Evaluated on a diverse set of SEC filings and real analyst tasks, our system demonstrates significant improvements in accuracy, controllability, and scalability over traditional and monolithic LLM baselines.
2. Related Work
2.1. Financial Information Extraction
Extracting structured data from financial documents such as 10-K, 10-Q, and earnings call transcripts has been a long-standing challenge. Early approaches employed rule-based parsing and template-matching techniques to identify financial entities and events, but these methods tend to be brittle under formatting variation or narrative complexity (Sheikh and Conlon, 2012; Costantino and Coletti, 2008). Supervised event-based extraction systems have also been explored, such as Malik et al.’s classifier-based framework for quantitative financial events (Malik et al., 2011) and SENTiVENT, which provides fine-grained labeled corpora for economic news event extraction (Jacobs and Hoste, 2022). While these systems demonstrated strong performance within narrow domains, they require substantial feature engineering and lack generalizability across document types.
More recently, LLM-based extractors have been introduced to capture KPIs in a zero- or few-shot manner (Costantino and Morgan, 1997). While promising, these models often lack domain grounding and tend to hallucinate values or overlook subtle context such as fiscal guidance, non-GAAP metrics, or period ranges. In contrast, our system explicitly integrates domain-specific logic and validation steps into the extraction pipeline to ensure reliability and robustness across diverse financial disclosures.
2.2. LLMs and Financial Document Understanding
Large language models (LLMs) have been increasingly applied to financial document tasks such as summarization, question answering, and data extraction (Yang et al., 2024; Li et al., 2024a; Shah et al., 2024; Ziegler, 2024; Srivastava et al., 2024). While they demonstrate promising capabilities in processing general text, recent studies highlight fundamental limitations when applied to semi-structured or domain-specific financial content.
For example, Srivastava et al. (2024) find that LLMs struggle with arithmetic reasoning in financial QA, particularly when aggregating values across tables or interpreting fiscal context. Similarly, Ziegler (Ziegler, 2024) observes that LLMs lack consistent understanding of financial and environmental terminology, often requiring external schema knowledge to extract accurate insights. Shah et al. (Shah et al., 2024) further note that despite their potential, current models fail to maintain consistency across multi-document financial QA, particularly when referencing nuanced company disclosures.
Moreover, hallucination, factual drift, and inability to trace predictions to source spans remain persistent issues (Li et al., 2024a). These challenges are compounded when LLMs are applied in a single-pass, end-to-end setting, where intermediate errors cannot be audited or corrected.
Our system addresses these limitations by separating responsibilities across two dedicated agents. The Extraction Agent ensures grounded, validated KPI outputs using financial logic, while the Text-to-SQL Agent enables structured querying over these results. This modularity allows iterative refinement and greater reliability, addressing key shortcomings identified in recent LLM-based financial IR studies.
2.3. Text-to-SQL and Financial Querying
Text-to-SQL systems aim to translate natural language questions into executable SQL over structured data. While general-purpose models have advanced considerably with benchmarks like Spider (Singh et al., 2025), their effectiveness in financial domains remains limited. Challenges arise from ambiguous query phrasing, implicit units or timeframes, and domain-specific schema structures (Wretblad and Riseby, 2024; Zhang et al., 2024).
To address this, recent studies have introduced financial-specific Text-to-SQL datasets such as FinSQL (Zhang et al., 2024) and BookSQL (Kumar et al., 2024), which provide training corpora grounded in accounting or financial reporting. However, even with domain adaptation, models often produce syntactically valid but semantically incorrect queries-such as mixing YoY and QoQ values, misidentifying column types, or violating fiscal constraints (Song et al., 2024; Wretblad and Riseby, 2024).
Our framework mitigates these issues by embedding schema-aware prompting, constraint validation, and semantic consistency checks into a dedicated Text-to-SQL Agent. This modular design allows fine-grained control over unit interpretation, time alignment, and execution correctness-enabling robust, analyst-friendly query interfaces for dynamic financial datasets.
2.4. Multi-Agent Architectures for Modular Reasoning
Recent advances in LLM-based multi-agent systems have shown that task decomposition and agent coordination can significantly improve reasoning robustness, modularity, and scalability (Tran et al., 2025; Li et al., 2024b; Agashe et al., 2023). Rather than relying on a single-shot monolithic LLM pipeline, multi-agent frameworks assign subtasks to specialized agents with explicit responsibilities and interaction protocols. This structure allows intermediate outputs to be validated, debugged, and iteratively improved.
For example, AgentNet (Yang et al., 2025) explores decentralized architectures for coordinating LLM agents in complex planning settings, while LLM-Coordination (Agashe et al., 2023) analyzes how cognitive structures such as Theory of Mind can enhance collaborative task completion. Surveys by Li et al. (Li et al., 2024b) and Tran et al. (Tran et al., 2025) emphasize that multi-agent designs offer benefits such as reduced hallucination, explicit memory sharing, and better long-range task alignment.
Figure 1. Comparison between the conventional manual pipeline and our proposed agentic system for financial KPI extraction and retrieval. While the conventional approach (Top) relies on labor-intensive human processes across both extraction and retrieval steps, our method (Bottom) automates the workflow using two agents—Extraction Agent and Text-to-SQL Agent—which collaborate to deliver scalable, accurate, and end-to-end structuring and querying of financial data.
[…]
Diagram comparing manual and agentic systems for financial KPI extraction and retrieval. The manual system involves human steps, while the agentic system uses two collaborating agents for automation.
3. Method
This section introduces our agentic system for automated financial KPI extraction and querying. We begin by outlining the challenges of manual workflows that analysts face when working with financial documents. We then present a modular two-agent architecture consisting of (1) an Extraction Agent, which transforms raw filings into structured KPI records, and (2) a Text-to-SQL Agent, which enables natural language querying over these records. Finally, we describe the technical architecture that supports scalable deployment and integration.
3.1. Motivation: Manual KPI Extraction is Inefficient
Financial analysts frequently extract key performance indicators (KPIs) from documents like 10-Ks, 10-Qs, earnings releases, and investor transcripts. As illustrated in Fig. 2, this process involves manually identifying relevant statements, interpreting financial terms, normalizing figures (e.g., converting ”22–24%” to midpoint 23%), and transferring them to downstream tools like spreadsheets or databases. This pipeline is slow, error-prone, and often lacks standardization—especially when dealing with ranges, units, and ambiguous time periods.
Figure 2. Manual KPI extraction process from earnings releases (e.g., Chipotle). Analysts must parse text, resolve numeric ranges, and enter structured data manually.
3.2. System Overview: A Two-Agent Architecture
To address these inefficiencies, we propose a modular system based on a two-step multi-agent architecture, shown in Fig. LABEL:fig:system_architecture. The system separates the extraction and querying components into specialized agents, each designed to optimize for financial accuracy, controllability, and interpretability.
The pipeline proceeds as follows:
  1. (1)
    Document Saving: Raw inputs including PDFs, HTML earnings pages, and transcripts are stored.
  2. (2)
    Layout & Parsing: Documents are OCR-processed (if necessary) and segmented by logical sections.
  3. (3)
    Preprocessing: Text is normalized; tables and numeric spans are detected.
  4. (4)
    Embedding & Structuring: Domain-specific embedding models identify KPI-relevant segments.
  5. (5)
    Step 1 – Extraction Agent: KPIs, fiscal periods, and metadata are extracted and validated.
  6. (6)
    Step 2 – Text-to-SQL Agent: Analyst queries are translated into executable SQL via a schema-aware model.
3.3. Extraction Agent
The Extraction Agent transforms raw financial documents into structured, machine-readable KPI records through a seven-stage process combining neural reasoning and rule-based logic. Fig. 3 shows performance across each step.
Figure 3. Evaluation breakdown for the Extraction Agent across its seven iterative stages. Each step contributes to higher accuracy, schema alignment, and unit/period consistency.
  • Raw Data Parsing: Layout-aware parsers segment filings into logical sections. OCR is applied where necessary.
  • Entity and Metric Detection: Domain-tuned prompting identifies candidate KPI spans using span-level and taxonomy-aware templates.
  • Contextual Span Extraction: Surrounding context like time periods and qualifiers (e.g., “guidance”, “non-GAAP”) is extracted alongside numeric values.
  • Preliminary Structuring: KPI spans are mapped into {metric, value, unit, period, qualifier} schema.
  • Domain Rule Injection: Heuristics normalize ranges, resolve units, disambiguate fiscal periods, and handle adjusted metrics.
  • QA-Driven Validation: Consistency is verified via targeted questions; inconsistent outputs are flagged or corrected.
  • Final Structuring and Output: Records are emitted as JSON with document metadata and confidence scores.
3.4. Text-to-SQL Agent
The Text-to-SQL Agent enables analysts to pose natural language questions over the structured KPI database. It translates queries like “What was the Q4 operating margin?” into executable SQL, handling financial semantics and schema variation.
  • Query Understanding: Parse target metrics, time constraints, aggregation logic, and qualifiers.
  • Schema Augmentation: Inject prompt context with schema metadata, column types, aliases, and sample rows.
  • SQL Generation: Use LLMs (e.g., Claude 3.7, GPT-4) with financial prompting and few-shot examples.
  • Constraint Validation: Check syntax, unit consistency, temporal alignment, and qualifier accuracy.
  • Execution Feedback: Test the query and rerank or regenerate if the output is implausible or empty.
  • Explanation Generation: Translate SQL into natural language summaries to support interpretability.
Figure 4. Iterative pipeline of the Text-to-SQL Agent. Each stage—from intent parsing and schema enrichment to constraint validation and execution testing—is designed to ensure high-fidelity SQL generation and interpretability.
3.5. Overall Pipeline
The overall pipeline is implemented on Databricks and supports scalable, auditable financial document processing.
  • Document Ingestion: Secure API/file-drop ingestion of filings into a unified data lake.
  • Parsing and Preprocessing: Layout-aware parsing and normalization via PDF/HTML/OCR pipelines.
  • LLM Orchestration: Claude 3.7 is used for structured extraction through prompt-based workflows.
  • Data Storage: Validated KPI records are stored in Delta Lake tables with schema guarantees.
  • Text-to-SQL Interface: Analysts query data using natural language interfaces mapped to the current schema.
  • Monitoring: Rule-based and MLFlow-based tracking ensures reliability, version control, and auditability.
This architecture supports extensibility across new document types, robustness to schema drift, and integration with third-party systems or human-in-the-loop workflows.
4. Experiments
4.1. Experimental Setup
To evaluate the effectiveness of our multi-agent framework, we simulate real-world analyst workflows involving the extraction of KPIs from SEC filings (10-Ks, 8-Ks, earnings releases) and natural language querying over those KPIs. Our evaluation uses:
  • Documents: 1,200 SEC filings from publicly traded companies.
  • Queries: 312 natural language questions derived from hedge fund analyst logs.
We compare three configurations:
  • Baseline A (GPT-4 Zero-Shot): Direct prompting without schema or constraints.
  • Baseline B (Claude 3.7 + Domain Prompt): Structured prompting with schema knowledge.
  • Ours (Claude 3.7 + Rule Injection): Full pipeline with rule-based validation and semantic constraints.
Table 1. Performance on individual agent: Extraction Agent
Metric Value
Metric Precision 95.3%
Structuring Accuracy 91.6%
Unit Error Rate 2.3%
QA Match Rate 92.3%
Schema Compliance 100.0%
Table 2. Performance on individual agent: Text-to-SQL Agent
Metric Value
Intent Classification Accuracy 95.6%
SQL Syntax Validity 98.2%
Semantic Constraint Pass Rate 91.3%
Execution Success Rate 96.4%
Top-1 SQL Accuracy (Human Eval) 91.2%
Explanation Relevance 90.5%
4.2. Component-Level Evaluation and End-to-End Pipeline Comparison
Table 3. Comparison of overall pipeline (Extraction Agent & Text-to-SQL Agent) across configurations
Config. Extraction F1 Query Accuracy Unit Error Rate Throughput Gain
GPT-4 (naive zero-shot) 72.4% 81.3% 18.5% 1.0×
Claude 3.7 (domain prompt) 84.7% 89.6% 7.2% 1.5×
Claude + Rule Injection 91.2% 94.8% 2.3% 2.4×
We first assess each agent in isolation to measure its contribution to overall accuracy in Table 1 and  2. Next, we evaluate the combined system across the three configurations in Table 3. The results show that structured prompting and rule-based validation significantly improve both extraction and query accuracy while reducing error rates and increasing throughput.
Prompt Engineering Impact.
Domain-specific prompt tuning (Claude 3.7) improves extraction F1 by +12.3 points over GPT-4 and reduces unit error rate by 60%, demonstrating the value of schema-aware prompting.
Rule Injection Effect.
Adding validation and heuristics significantly reduces semantic and numeric misalignment, especially for period mismatches and GAAP vs. non-GAAP misinterpretation.
Query Reliability.
Constraint-validated SQL generation yields a +13.5 point gain in query execution accuracy compared to zero-shot baselines, reflecting better alignment between user intent and schema-constrained SQL.
System Efficiency.
Our full framework achieves a 2.4× throughput gain by minimizing error propagation and reducing manual normalization overhead.
4.3. Error Case Examples
  • GPT-4 Zero-Shot: Interpreted “operating margin guidance of 15–17%” as actual Q4 value (16.0%).
  • Claude (no rule): Extracted 16.0% correctly but failed to flag it as guidance.
  • Claude + Rules: Identified range as guidance and excluded it from GAAP-specific queries.
Our modular, rule-enhanced multi-agent architecture enables robust, interpretable financial QA. By decomposing the pipeline into extraction, validation, and SQL generation, we gain control over accuracy, explainability, and system-level efficiency. These improvements are particularly valuable for real-world analyst workflows, where numeric precision and auditability are paramount.
5. Conclusion
We presented a modular, multi-agent framework for extracting and querying structured information from unstructured financial documents such as 10-Ks, 8-Ks, and earnings transcripts. By decomposing the pipeline into a dedicated Extraction Agent and a Text-to-SQL Agent, our system addresses key challenges in financial information retrieval-including document layout variation, unit normalization, and semantic ambiguity in natural language queries.
Through prompt engineering and domain-specific validation, our architecture outperforms baseline LLM setups in both accuracy and robustness, enabling analysts to access high-fidelity KPI data with significantly reduced manual effort. The framework’s separation of concerns also allows for targeted tuning, iterative refinement, and easier auditability.
Future work includes expanding support to multilingual filings, integrating richer temporal logic, and enabling cross-document reasoning over longer time horizons. We believe this approach represents a scalable, trustworthy foundation for deploying generative AI in financial research workflows.
Appendix A Prompt Engineering Examples
This section compares two prompting strategies used in the Extraction Agent: a naive zero-shot prompt and a domain-aware prompt that incorporates schema, unit normalization, and period anchoring.
Table 4. Prompt Engineering Comparison: Zero-Shot vs Domain-Aware
Zero-Shot Prompt Domain-Aware Prompt
Prompt
Extract key financial figures from the text below.
Prompt
Extract financial KPIs from the earnings text using the schema below. Output in JSON. Normalize units and link metrics to fiscal periods.
Text
In Q1 2024, revenue grew 12% YoY to \$4.3 billion, beating consensus by \$150 million.
Text
Same as left. Schema includes Period, Revenue (B), Revenue_YoY_Growth, and Consensus_Delta (M).
Output (Expected)
{ Revenue: "4.3B", Growth: "12%", Period: "Q1 2024" }
Output (Expected)
{ "Period": "Q1 2024", "Revenue": {value: 4.3, unit: "B"}, "Revenue_YoY_Growth": "12%", "Consensus_Delta": {value: 150, unit: "M"} }
Notes
Lacks structure; prone to inconsistent formatting and unit mismatches.
Notes
Improves consistency, reduces hallucinations, enables downstream parsing and SQL conversion.
Appendix B Ablation Study: Rule-based Validation
As shown in Table 5, disabling rule-based validation leads to significantly higher unit and period errors.
Table 5. Ablation: Without Rule-Based Validation
Metric With Rules Without Rules $\delta$ Error
Extraction Precision 91.2% 82.6% -8.6%
Unit Error Rate 2.3% 12.9% +10.6%
Period Misalignment Rate 3.1% 15.4% +12.3%
QA Match Accuracy 90.4% 78.7% -11.7%
SQL Validity Pass Rate 94.8% 86.0% -8.8%
Prompt A: Naive Zero-Shot Prompt (No Agent Iteration)
Prompt:
”Extract all financial metrics mentioned in the following text and return them as a JSON object.”
Example Input Text:
”Operating margin in Q4 2024 was 14.6%, up from 14.4% last year. Revenue grew 15.2% to \$2.52 billion. The company expects operating margin to be between 15–17% in FY 2025.”
LLM Output: “‘json ”Operating margin”: ”14.6”, ”Revenue”: ”\$2.52 billion”

### v2

@@
 Abstract.
-Extracting structured and quantitative insights from unstructured financial filings is essential in investment research, yet remains time-consuming and resource-intensive.
+Extracting and retrieving structured and quantitative insights from unstructured financial filings is essential in investment research, yet remains time-consuming and resource-intensive.
 Conventional approaches in practice rely heavily on labor-intensive manual processes, limiting scalability and delaying the research workflow.
-In this paper, we propose an efficient and scalable method for accurately extracting quantitative insights from unstructured financial documents, leveraging a multi-agent system composed of large language models.
+In this paper, we propose an efficient and scalable method for both extracting and retrieving financial insights accurately, leveraging a multi-agent system composed of large language models.
 Our proposed multi-agent system consists of two specialized agents: the Extraction Agent and the Text-to-SQL Agent.
@@
 1. Introduction
-The financial industry generates vast volumes of unstructured documents-10-Ks, 10-Qs, 8-Ks, DEF14As, and investor memos-that are central to investment research and decision-making.
-Analysts have traditionally extracted key performance indicators (KPIs) from these documents by manually scanning text, copying metrics, normalizing formats, and writing internal summaries.
-This process is not only time-consuming but also error-prone and unscalable in modern research environments where timeliness and breadth of coverage are essential.
-While recent large language models (LLMs) have shown promise in structuring unstructured data, their application in financial workflows remains unreliable.
-In the extraction phase, LLMs often misinterpret numeric guidance, extract semantically incorrect values (e.g., YoY instead of QoQ), or omit critical metrics entirely.
-These issues stem from a lack of domain-specific reasoning, layout ambiguity, and difficulty distinguishing factual signals from narrative commentary.
-Even when extraction is successful, querying this data via natural language introduces further challenges: text-to-SQL models are sensitive to phrasing, schema variation, and inherent LLM randomness, often leading to invalid or misleading responses.
-These problems are compounded by the fact that most LLM-based systems rely on monolithic, single-shot inference pipelines that lack transparency, fail to isolate errors, and cannot iteratively refine intermediate outputs.
-This makes it difficult to enforce consistency, inject domain rules, or apply structured validation across stages.
-To address these challenges, we propose a scalable, document-agnostic multi-agent framework for financial document understanding.
+Investment research depends critically on the analysis of unstructured corporate disclosures.
+In practice, analysts often extract key performance indicators (KPIs) either by manually scanning lengthy documents or by locating figures scattered across external sources and internal summaries.
+As firms continue to release a growing volume of earnings transcripts, financial reports, and regulatory filings, such manual processes have become increasingly time-consuming and impractical.
+To ensure both timeliness and breadth of coverage, scalable methods for extracting and retrieving insights from these texts are now essential to informed decision-making.
+Recent advances in large language models (LLMs) have opened new possibilities for processing unstructured documents at scale, offering a promising foundation for automating financial information extraction and retrieval.
+These models can, in principle, transform complex textual data into structured insights and support natural language interactions with document content.
+However, despite their potential, LLMs yet remain unreliable in real-world financial workflows.
+When used in isolation, they often misinterpret numeric guidance, extract semantically incorrect values, or omit critical metrics altogether.
+Such failures are largely due to limitations in layout sensitivity and the difficulty of distinguishing factual signals from surrounding narrative.
+To address these shortcomings, recent work has explored multi-agent frameworks, in which multiple LLMs collaborate to validate, cross-check, and refine each other’s outputs—improving accuracy, reducing hallucinations, and enhancing reliability in financial document analysis.
+In this context, we propose an efficient and scalable method for both extracting and retrieving financial insights accurately, leveraging a multi-agent system composed of LLMs.
 Our system consists of two specialized agents: an Extraction Agent, which parses and validates KPIs using domain-tuned prompts and logic, and a Text-to-SQL Agent, which translates analyst queries into executable SQL over the structured output.
@@
     Step 2 – Text-to-SQL Agent: Analyst queries are translated into executable SQL via a schema-aware model.
+Figure 3.
+Overview of our multi-agent system for financial document understanding.
+The Extraction Agent processes unstructured documents to extract structured KPIs using LLMs (LLaMA, Claude) and stores them in Delta Lake via Apache Spark.
+The Text-to-SQL Agent enables natural language querying by generating metadata-aware SQL queries, retrieving results through Databricks SQL.
+The system is built on the Databricks stack, including MLflow for auditing and AI Gateway for traceability.
 3.3. Extraction Agent
 The Extraction Agent transforms raw financial documents into structured, machine-readable KPI records through a seven-stage process combining neural reasoning and rule-based logic.
-Fig.
-3 shows performance across each step.
-Figure 3. Evaluation breakdown for the Extraction Agent across its seven iterative stages. Each step contributes to higher accuracy, schema alignment, and unit/period consistency.
+The top side of Fig.
+3 shows its detailed steps.
   • Raw Data Parsing: Layout-aware parsers segment filings into logical sections. OCR is applied where necessary.
@@
 It translates queries like “What was the Q4 operating margin?” into executable SQL, handling financial semantics and schema variation.
+The bottom side of Fig.
+3 shows its detailed steps.
   • Query Understanding: Parse target metrics, time constraints, aggregation logic, and qualifiers.
@@
   • Explanation Generation: Translate SQL into natural language summaries to support interpretability.
-Figure 4.
-Iterative pipeline of the Text-to-SQL Agent.
-Each stage—from intent parsing and schema enrichment to constraint validation and execution testing—is designed to ensure high-fidelity SQL generation and interpretability.
 3.5. Overall Pipeline
@@
 4. Experiments
+Table 1. Comparison of overall pipeline (Extraction Agent & Text-to-SQL Agent) across configurations
+Config. Extraction F1 Query Accuracy Unit Error Rate Throughput Gain
+GPT-4o (naive zero-shot) 72.4% 81.3% 18.5% 1.0×
+Claude-3.7-Sonnet (domain prompt) 84.7% 89.6% 7.2% 1.5×
+Claude-3.7-Sonnet + Rule Injection 91.2% 94.8% 2.3% 2.4×
 4.1. Experimental Setup
@@
   • Queries: 312 natural language questions derived from hedge fund analyst logs.
+Each document was manually annotated by trained financial analysts to identify ground-truth KPIs, including revenue, operating income, free cash flow, and forward-looking guidance.
+Annotations were cross-validated for consistency and aligned with GAAP definitions where applicable.
 We compare three configurations:
@@
 Schema Compliance 100.0%
-Table 2. Performance on individual agent: Text-to-SQL Agent
-Metric Value
-Intent Classification Accuracy 95.6%
-SQL Syntax Validity 98.2%
-Semantic Constraint Pass Rate 91.3%
-Execution Success Rate 96.4%
-Top-1 SQL Accuracy (Human Eval) 91.2%
-Explanation Relevance 90.5%
 4.2. Component-Level Evaluation and End-to-End Pipeline Comparison
-Table 3. Comparison of overall pipeline (Extraction Agent & Text-to-SQL Agent) across configurations
-Config. Extraction F1 Query Accuracy Unit Error Rate Throughput Gain
-GPT-4 (naive zero-shot) 72.4% 81.3% 18.5% 1.0×
-Claude 3.7 (domain prompt) 84.7% 89.6% 7.2% 1.5×
-Claude + Rule Injection 91.2% 94.8% 2.3% 2.4×
-We first assess each agent in isolation to measure its contribution to overall accuracy in Table 1 and  2.
+We first assess each agent in isolation to measure its contribution to overall accuracy in Table 2 and  3.
 Next, we evaluate the combined system across the three configurations in Table 1.
@@
 Our full framework achieves a 2.4× throughput gain by minimizing error propagation and reducing manual normalization overhead.
+Table 3. Performance on individual agent: Text-to-SQL Agent
+Metric Value
+Intent Classification Accuracy 95.6%
+SQL Syntax Validity 98.2%
+Semantic Constraint Pass Rate 91.3%
+Execution Success Rate 96.4%
+Top-1 SQL Accuracy (Human Eval) 91.2%
+Explanation Relevance 90.5%
 4.3. Error Case Examples